Circular No. 6428

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 N1 A. Hale, Cloudcroft, NM, reports the visual discovery of a comet by

Howard Brewington (also of Cloudcroft). Available observations:

1996 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 4.177  11^{h}33.^{m}5  +4^{o}06′  10 Brewington
4.35560 11 34 01.^{s}33 +4 13 13′′.7  Garradd
4.35647 11 34 01.49 +4 13 16.8  _{”}
4.35739 11 34 01.62 +4 13 20.9  _{”}
4.36856 11 34 03.61 +4 13 59.6  _{”}
4.36955 11 34 03.73 +4 14 02.9  _{”}

H. J. Brewington (Cloudcroft, NM). 0.20-m reflector. G. J. Garradd (Loomberah, N.S.W.). 0.25-m f/4.1 reflector + CCD. Some

cirrus clouds. Coma diameter ∼ 5′ on a 300-s exposure.

4U 1636–536 AND 4U 0614+091 M. van der Klis, Astronomical Institute “Anton Pannekoek”, Univer-

sity of Amsterdam; J. van Paradijs, University of Amsterdam and University of Alabama in Huntsville; W. H. G. Lewin, Massachusetts Institute of Technology; F. K. Lamb, University of Illinois, Urbana/Champaign; B. Vaughan, California Institute of Technology; E. Kuulkers, Astrophysics Division, ESTEC, European Space Agency; and T. Augusteijn, European Southern Observatory, report: “RXTE observations of 4U 1636–536 performed on Feb. 28.24–28.35 UT show two simultaneous kHz quasiperiodic oscillations (QPOs) at frequencies of 898.7 ± 1.4 and 1171 ± 11 Hz. Power spectral peak widths are 28 ± 5 and 88 ± 26 Hz (FWHM), respectively, and root-mean-square amplitudes 6.6 ± 0.2 and 6.1 ± 0.7 percent (7–20 keV). The higher-frequency peak is the fastest x-ray QPO yet reported. Two RXTE observations of 4U 0614+091 (cf. IAUC 6426) that were performed on Mar. 16.35–16.46 and Apr. 13.41–13.54 each show a single kHz QPO peak, with frequencies, FWHMs, and rms amplitudes of 727 ± 5 Hz, 137 ± 18 Hz, and 15.8 ± 0.8 percent; and 630 ± 11 Hz, 126 ± 27 Hz, and 16.5 ± 1.5 percent (2–20 keV), respectively.”

NOVA CASSIOPEIAE 1995 Visual magnitude estimates (cf. IAUC 6358) by M. V. Zanotta, Milan:

Apr. 6.85 UT, 9.4; 15.12, 8.5; May 20.87, 9.5; June 10.95, 9.9; 30.06, 10.1.

1996 July 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green